# Inconsistency with De Sitter Spacetime in a New Approach to Gravitational Particle Production

### v2

@@ Abstract
We study the recent Physical Review Letter Wondrak:2023zdi which presents a new mechanism for particle production and black hole evaporation through a spatially dependent temperature. This new temperature is comparable to the Hawking result near the black hole, but is very small far away, and therefore could be a small correction. Here we apply the proposed reasoning to the case of de Sitter space, finding that it over predicts the de Sitter temperature of a minimally coupled scalar by factor of $\approx 4.3$ and over predicts the particle production rate by a factor of $\approx 52$. For non-minimally coupled scalars, it has other various problems; it predicts a negative particle production for conformal, or nearly conformal, coupled scalars; it predicts unsuppressed productions of heavy scalars. This all demonstrates an inconsistency in the proposed formalism.
@@
where $\Omega_{\mu\nu}$ is the gauge-field curvature $\Omega_{\mu\nu}=iqF_{\mu\nu}$. By utilizing the effective action for scalar particles in curved space time, they note that the combination $\Omega_{\mu\nu}\Omega^{\mu\nu}/2$ appears in a way analogous to several contractions of Riemann, which they identify. Through this analogy, they suggest that a gravitational field should also produce pairs of particles with characteristic energy
@@
where $\xi$ is the non-minimal coupling parameter of the scalar (defined through $-\xi R\phi^{2}/2$ in the action); $\xi=0$ corresponds to minimally coupled and $\xi=1/6$ corresponds to conformally coupled. Ref. Wondrak:2023zdi focuses on the vacuum case (Schwarzschild black holes) in which $R_{\mu\nu}=0$ and $R=0$, which simplifies this expression; but we shall track the full expression here.
@@
To apply this to de Sitter space, we integrate over the Hubble volume Vol $=4\pi/(3H^{3})$ and again use the values of $R_{\mu\nu\alpha\beta}$ for de Sitter to obtain
@@
Note added.— While we wrote our paper, another work appeared in Ref. Ferreiro:2023jfs which also finds inconsistencies in the proposal. They examine the electromagnetism in more detail and comment on FLRW, while we examine de Sitter spacetime and the subtle dependence on $\xi$ in more detail (its complexity in conformal case, its precise value for minimally coupled, etc).

### v3

@@ Abstract
We study a claimed new mechanism for particle production and black hole evaporation through a spatially dependent temperature. This new temperature is comparable to the Hawking result near the black hole, but is very small far away, and therefore could be a small correction. Here we apply the proposed reasoning to the case of de Sitter space, finding that it over predicts the de Sitter temperature of a minimally coupled scalar by factor of $\approx 4.3$ and over predicts the particle production rate by a factor of $\approx 52$. For non-minimally coupled scalars, it has other various problems; it predicts a negative particle production for conformal, or nearly conformal, coupled scalars; it predicts unsuppressed productions of heavy scalars. This all demonstrates an inconsistency in the proposed formalism.
@@
where $\Omega_{\mu\nu}$ is the gauge-field curvature $\Omega_{\mu\nu}=iqF_{\mu\nu}$. By utilizing the effective action for scalar particles in curved space time, they note that the combination $\Omega_{\mu\nu}\Omega^{\mu\nu}/2$ appears in a way analogous to several contractions of Riemann, which they identify. Through this analogy, and using an effective action $\mbox{Im}[W]_{\rm eff}={1\over 384\pi}\int d^{4}x\sqrt{-g}\,E_{\mbox{\tiny{gravity}}}^{4}$, they suggest that a gravitational field should also produce pairs of particles with characteristic energy
@@
where $\xi$ is the non-minimal coupling parameter for a scalar (defined through $\xi R\phi^{2}/2$ in the action); $\xi=0$ corresponds to minimally coupled and $\xi=1/6$ corresponds to conformally coupled. Ref. Wondrak:2023zdi focuses on the vacuum case (Schwarzschild black holes) in which $R_{\mu\nu}=0$ and $R=0$, which simplifies this expression; but we shall track the full expression here.
@@
To apply this to de Sitter space, we integrate over the Hubble volume Vol $4\pi/(3H^{3})$ and again use the values of $R_{\mu\nu\alpha\beta}$ for de Sitter to obtain
@@
Note added.— While we wrote our paper, another work appeared in Ref. Ferreiro:2023jfs which also finds inconsistencies in the proposal. They examine the case of electromagnetism in more detail and comment on FLRW, while we examine de Sitter spacetime and the subtle dependence on $\xi$ in more detail (its complexity in conformal case, its precise value for minimally coupled, etc).
In the original version of Ref. Ferreiro:2023jfs , it is suggested that there is particle production in generic FLRW spacetimes for the conformally coupled scalar. However this claim has 2 shortcomings: (i) The proposal of Ref. Wondrak:2023zdi cannot obviously be directly used for time dependent spacetimes, which includes generic FLRW spacetimes. This is why in our work we have focussed on one of the few nontrivial cosmological spacetimes which is stationary, namely de Sitter. (ii) For the conformally coupled scalar, there is not in fact particle production, as we showed here. Instead, as we first showed, for $\xi=1/6$ one instead has $E_{\mbox{\tiny{gravity}}}^{4}$ (or $T_{\mbox{\tiny{eff}}}^{4}$) being negative, corresponding to a complex value when taking the fourth root. It is only for $\xi$ outside of a certain window (see Figure 1) that the fourth root is real and so there is in fact particle production.
Later, after our first version appeared, Ref. Ferreiro:2023jfs put out a new version which included an appendix. In this appendix, they realized these two shortcomings in their original version and corrected them: (i) they mentioned the special case of de Sitter space and (ii) the fact that the production is negative for the conformally coupled scalar. (To show this, they found $\mbox{Im}[W]_{\rm eff}=-{12\over 192\pi}\int d^{4}x\sqrt{-g}\,H^{4}$. While the minus sign is correct, which follows from the complex values for $E_{\mbox{\tiny{gravity}}}$ we originally discovered, the prefactor of $12/(192\pi)$ is incorrect; the correct value is $1/(960\pi)$, which can be deduced from our above value for $E_{\mbox{\tiny{gravity}}}$ or $T_{\mbox{\tiny{eff}}}$.)
We summarize some of the central new points in our work: (i) the novelty of using de Sitter space (due to its stationarity), (ii) the precise computation for the action and the particle rates as a function of $\xi$, including the precise identification of positive versus negative production, or real versus complex effective energy, (iii) the identification that at large $\xi$ the approach would lead to a significant rate for an effectively heavy scalar in de Sitter space, rather than exponential suppression.